# Information Transport in Classic-Quantum Hybrid System

### v1

@@ V Two-Replica Hybridization
Each qubit replica (green) couples to its local environment through the interaction Hamiltonian $H_{RQ}$. At second order, two crosses denote the two interactions $H_{RQ}(t_{1})$ and $H_{RQ}(t_{2})$ between the qubit $Q$ and the reservoir R. Mixed terms involving a single $RQ$ only in a reservoir vanish. Comprehensive derivations of these diagrammatic rules—and their efficient resummations—can be found in Refs. [22, 24, 31].
@@
—Tracing out the main reservoir $e$ within each replica yields the canonical contribution [56]:
@@
with $\Delta^{(2)}_{e}$ denotes the Lamb shift induced by the environment and defined as follows $\Delta_{e}=\Pi_{01,10}^{(0,1)}(\omega)-\Pi_{10,01}^{(0,1)}(-\omega)$. The transition rates $\Gamma_{e\uparrow}=\Gamma_{e\downarrow}\,\exp({-\omega/T_{e}})=\chi_{{}_{e}}\,n_{B}({\omega}{T_{e}})$ and $\chi_{{}_{e}}$ its dynamical susceptibility of the environment. The definition of $\Pi_{01,10}^{(0,1)}$ follows the generalized definition of correlators and the generalized KMS theorem on $M$ replicas, discussed in details and derived diagrammatically within the KAN formalism [22].
@@
Figure 6: Entropy flow versus drive amplitude. Numerically computed probe Rényi entropy current $\dot{S}_{\alpha}$ plotted against the reduced drive $\Omega/\Gamma_{b}$ for different choices of $\omega/T$. The results in (a) and (c) follow the hybridization expression of Eq. (14); while (b) and (d) show the standard weak-coupling result of Eq. (3). Panels (a) and (b) correspond to strong probe coupling ($\Gamma_{b}=\Gamma_{e}$), while (c) and (d) illustrate the weak-coupling limit ($\Gamma_{b}/\Gamma_{e}\ll 1$).
@@
—Summing the diagrams of Fig. (4) and the integration over the internal time coordinate yields the following Liouvillian matrix:
@@
Such cross terms have ben depicted in diagrams of Fig. (5) in which each replica contains only one interaction in time, when is summed over for all possible such exchanges will produce
@@
Collecting all terms yields the total Liouvillian matrix, a compact $16\times 16$ matrix that governs the dynamics of the replicated density operator $\hat{R}_{2}$ and, via the replica limit, directly produces the Rényi-2 entropy production rate. The total Liouvillian superoperator for $M=2$ can been explicitly seen in Appendix (B). Throughout the unindex $\Gamma$ is used sometime instead of $\Gamma_{e}$.
After determining the Liouvillian matrix, we construct an orthonormal supervector basis $|\Phi\rangle\!\rangle$ that diagonalizes the total Liouvillian superoperator. The entropy flow of Eq. (14) is the smallest Liouvillian eigenvalue.
@@
To examine the strong-coupling regime $\Gamma_{b}=\Gamma$ in more detail, we directly compare the two formalisms by plotting them together in Fig. 7. The figure shows the entropy-flow results from the hybridization formalism, $\Gamma^{(2)}_{\min}$ (red curves), and from the standard formalism, $\mathcal{F}_{2}$ (blue curves), as functions of $\Omega/T_{e}$ for various renormalized drive amplitudes $\Omega/\Gamma_{b}$. The red curves represent the hybridized results, while the blue curves correspond to the weak-coupling approach; in each color set, darker shades indicate stronger drive amplitudes.
@@ VI M-Replica Case
We now extend the replica construction from two copies to an arbitrary integer number $M$. For each extra replica, the generalized density operator becomes a tensor acting on a Hilbert space of dimension $2^{M}$, so its Liouvillian super-operator lives in a space of dimension $2^{2M}$. As long as $M$ is an integer, all of the required linear-algebraic objects—operators, tensor products, traces—are perfectly well-defined.
@@
Same-world superoperators.
@@
Figure 9: Von Neumann entropy for a hybridized system with $\Gamma_{b}=\Gamma_{e}$ as a function of $\omega/T_{e}$, shown for weak and strong driving amplitudes. The dashed line in blue color labeled as $F_{vN}$ represents the weak-coupling probe estimate from Eq. (5). Thinner dashed line labeled with $F_{vN}^{\rm{incoh}}$ denotes the incoherent von Neumann entropy flow without considering the cross-world contributions.
Figure 9 shows the von Neumann entropy as a function of $\omega/T_{e}$ for several driving amplitudes. The corresponding von Neumann entropy estimates obtained from the fitting procedure are also presented in Fig. 9 for weak and strong driving. The dashed line lablled with $S_{vN}$ represents the von Neumann entropy current evaluated directly from Eq. (5), which shows quantum coherent contribution in comparison to incoherent-only part of the flow presented by thinner dashed line labeled as $S_{vN}^{\rm{incoh}}$.
@@
In all curves, lowering the reservoir temperature suppresses the entropy, and in the case of the new hybridized entropy-flow estimation, this suppression follows an almost power-law behavior, particularly in the strong-driving limit, $\Omega\lesssim\Gamma_{e}$. The primary source of the observed deformation is that the calculation was performed for $\Gamma_{b}=\Gamma$, which makes the direct evaluation of entropy flow from Eq. (5) less accurate. The power-law dependence of the hybridized von Neumann entropy can be expressed as
$S_{vN}\propto\left(\frac{\omega}{T_{e}}\right)^{\eta},$ (29)
with $\eta\approx-1$. This indicates that, in the strong-coupling regime, lowering the temperature leads to a faster suppression of entropy compared to previous calculations based on other formalisms, with the effect being more pronounced at higher driving amplitudes. Similarly one can expect that the increase of external temperature helps to speed up in entropy production.
@@ Conclusion
The extension to an arbitrary number of replicas $M$ permits the estimation of the von Neumann entropy current by fitting the $M$-dependence of Rényi entropy flows. The extracted scaling follows a power-law suppression with temperature, $S_{\mathrm{vN}}\propto(\omega/T_{e})^{\eta}$, with $\eta\approx-1$. This reveals a more rapid entropy suppression in the strong-coupling regime than predicted by earlier formalisms, especially under strong driving conditions.

### v2

@@ V Two-Replica Hybridization
Each qubit replica (orange) couples to its local environment through the interaction Hamiltonian $H_{RQ}$. At second order, two crosses denote the two interactions $H_{RQ}(t_{1})$ and $H_{RQ}(t_{2})$ between the qubit $Q$ and the reservoir R. Mixed terms involving a single $RQ$ only in a reservoir vanish. Comprehensive derivations of these diagrammatic rules—and their efficient resummations—can be found in Refs. [22, 24, 31].
@@
—Tracing out the modes of reservoir $e$ within each replica yields the canonical contribution [63]:
@@
with $\Delta^{(2)}_{e}$ denotes the Lamb shift induced by the environment and defined as follows $\Delta_{e}=\Pi_{01,10}^{(0,1)}(\omega)-\Pi_{10,01}^{(0,1)}(-\omega)$. The transition rates $\Gamma_{e\uparrow}=\Gamma_{e\downarrow}\,\exp({-\omega/T_{e}})=\chi_{{}_{e}}\,n_{B}({\omega}{T_{e}})$ and $\chi_{{}_{e}}$ its dynamical susceptibility of the environment. The definition of $\Pi_{01,10}^{(0,1)}$ follows the generalized definition of correlators and the generalized KMS theorem on $M$ replica, discussed in details and derived diagrammatically within the KAN formalism [22].
@@
Figure 6: Entropy flow versus drive amplitude. Numerically computed probe Rényi entropy current $\dot{S}_{2}$ plotted against the reduced drive $\Omega/\Gamma_{b}$ for different choices of $\omega/T$. The results in (a) and (c) follow the hybridization expression of Eq. (13); while (b) and (d) show the standard weak-coupling result of Eq. (3). Panels (a) and (b) correspond to strong probe coupling ($\Gamma_{b}=\Gamma_{e}$), while (c) and (d) illustrate the weak-coupling limit ($\Gamma_{b}/\Gamma_{e}\ll 1$).
@@
—Summing the diagrams of Fig. (4) and the integration over the internal time coordinate yield the following Liuvillian matrix:
@@
Such cross terms have been depicted in diagrams of Fig. (5) in which each replica contains only one interaction in time, when is summed over for all possible such exchanges will produce
@@
Collecting all terms yields the total Liouvillian matrix, a compact $16\times 16$ matrix that governs the dynamics of the replicated density operator $\hat{R}_{2}$ and, via the replica limit, direct produces the Rényi-2 entropy production rate. The total Liouvillian superoperator for $M=2$ has been explicitly seen in Appendix (B).
After determining the Liouvillian matrix, we construct an orthonormal supervector basis $|\Phi\rangle\!\rangle\}$ that diagonalizes the total Liouvillian superoperator. The entropy flow obtained from Eq. (13) is the smallest Liouvillian eigenvalue.
@@
To examine the strong-coupling regime $\Gamma_{b}=\Gamma$ in more detail, we directly compare the two formalisms by plotting them together in Fig. (7). The figure shows the entropy-flow results from the hybridization formalism, $\Lambda^{(2)}_{\min}$ (red curves), and from the standard formalism, $\mathcal{F}_{2}$ (blue curves), as functions of $\omega/T_{e}$ for various renormalized drive amplitudes $\Omega/\Gamma_{b}$. The red curves represent the hybridized results, while the blue curves correspond to the weak-coupling approach; in each color set, darker shades indicate stronger drive amplitudes.
@@ VI M-Replica Case
We now extend the replica construction from two copies to an arbitrary integer number $M$. For each extra replica, the generalized density operator becomes a tensor acting on a Hilbert space of dimension $2^{M}$, so its Liouvillian super-operator acting on space of dinmension $2^{2M}$ and lives in a space of dimension $2^{4M}$. As long as $M$ is an integer, all of the required linear-algebraic objects—operators, tensor products, traces—are perfectly well-defined.
@@
Same-world super-operators.
@@
Figure 9: Von Neumann entropy flow derived from strong interaction method for a hybridized system with $\Gamma_{b}=\Gamma_{e}$ as a function of $\omega/T_{e}$, used to evaluate weak (blue dots) and strong (orange dots) driving amplitudes. The dashed line in blue dashed color labeled as $F_{vN}$ represents the weak-coupling probe estimate from Eq. (5). Thinner (green) dashed line labeled with $F_{vN}^{\rm{incoh}}$ denotes the incoherent von Neumann entropy flow without considering the cross-world contributions.
Figure 9 shows the von Neumann entropy as a function of $\omega/T_{e}$ for several driving amplitudes. The corresponding von Neumann entropy estimates obtained from the fitting procedure are also presented in Fig. 9 for weak and strong driving. The dashed line lablled with $F_{vN}$ represents the von Neumann entropy current evaluated directly from Eq. (5), which shows quantum coherent contribution in comparison to incoherent-only part of the flow presented by thinner dashed line labeled as $F_{vN}^{\rm{incoh}}$.
@@
In all curves, lowering the reservoir temperature suppresses the entropy, and in the case of the new hybridized entropy-flow estimation, this suppression follows an almost power-law behavior, particularly in the strong-driving limit. The primary source of the observed deformation is that the calculation was performed for $\Gamma_{b}=\Gamma$, which makes the direct evaluation of entropy flow from Eq. (5) less accurate. The power-law dependence of the hybridized von Neumann entropy can be expressed as
$S_{vN}\propto\left(\frac{\omega}{T_{e}}\right)^{\beta},$ (28)
with $\beta\approx-1$. This indicates that, in the strong-coupling regime, lowering the temperature leads to a faster suppression of entropy compared to previous calculations based on other formalisms, with the effect being more pronounced at higher driving amplitudes. Similarly one can expect that the increase of external temperature helps to speed up in entropy production.
@@ Conclusion
The extension to an arbitrary number of replicas $M$ permits the estimation of the von Neumann entropy current by fitting the $M$-dependence of Rényi entropy flows. The extracted scaling follows a power-law suppression with temperature, $S_{\mathrm{vN}}\propto(\omega/T_{e})^{\beta}$, with $\beta\approx-1$. This reveals a more rapid entropy suppression in the strong-coupling regime than predicted by earlier formalisms, especially under strong driving conditions.